# Timelike-bounded $dS_{4}$ holography from a solvable sector of the $T^{2}$ deformation

### v2

@@ 1 Introduction
The results Coleman:2021nor ; Batra:2024kjl were specific to a 3-dimensional bulk theory. In this note, we refine the method and point out that it readily generalizes to the case of four bulk dimensions, even though the special properties of the $T\bar{T}$ operator in two dimensions Zamolodchikov:2004ce ; Smirnov:2016lqw do not extend to the full $T^{2}$ operator in higher dimensions Hartman:2018tkw ; Taylor:2018xcy . For boundary dimension $d=3$ there is still a sector – containing the most entropic bands of energy levels – that is captured by a solvable deformation of the $AdS_{4}/CFT_{3}$ Hamiltonian similar to Coleman:2021nor . We define this solvable theory by restricting the $T^{2}$ deformation of AdS/CFT (as derived at large $N$ in Hartman:2018tkw and generalized here to include a constant $\Lambda_{3}$ addition at each step) to contain only diagonal components of the stress tensor and to take the form that it does in symmetric homogeneous states, depending only on energy density $E/V$ and pressure $-dE/dV$. The finite real spectrum of this Hamiltonian, manifest in equation (6) below, captures the Gibbons-Hawking entropy along with the radial geometry of the $dS_{4}$ cosmic horizon patch as a function of boundary size. The operator algebra, working at finite $N$, is type I, simpler than the type II result one obtains in the $G_{N}\to 0$ limit Chandrasekaran:2022cip .
@@ 1.1 Status of classical and quantum timelike boundaries
From the classical, low energy perspective it is well known that the initial boundary value problem is not well defined for broad classes of Dirichlet boundary conditions, at least linearized around simple backgrounds. For example, the radial constraint linearized around these backgrounds cannot be satisfied for an arbitrary fixed boundary geometry An:2021fcq . This has been straightforwardly proved for perturbations of a flat boundary in Minkowski spacetime (a flat ‘corner’) to nearby boundary metrics with nonzero intrinsic scalar curvature $\delta{\mathcal{R}}$: with vanishing unperturbed extrinsic curvature, the linearized perturbation to the radial constraint simply reads $\delta{\mathcal{R}}=0$. (This argument does not exclude solutions of the radial constraint with nontrivial extrinsic curvature scaling like $K\sim\sqrt{\delta g}$, which exists in many cases. Still, according to an argument in terms of Sobolev spaces in Anderson:2010ph these nonlinear solutions do not exist for generic fixed boundary metrics, as was seen in early works on near horizon physics Bredberg:2011xw ; Anninos:2011zn .) At least for flat corners at the linearized level there is also a violation of geometric uniqueness Anninos:2022ujl ; An:2021fcq : given initial data, it does not evolve to a unique solution at later times. At least for the Euclidean case, flat corners also lead to problems defining a propagator for perturbation theory Anderson:2006lqb ; Witten:2018lgb .
@@ 2 A solvable $T^{2}+\Lambda_{3}$ deformation at finite N
In the prescription Coleman:2021nor , one starts with $\eta=1$, deforms until $y$ reaches $y_{0}=3/c\pi^{2}$, interpolates to $\eta=-1$ (resolved into an uplift achieved by appropriate matter Dong:2010pm in Batra:2024kjl ), and proceeds back toward larger $y$ with the $\Lambda_{2}$ terms in place. This by itself accounts for the refined Gibbons-Hawking entropy and radial geometry of the bounded $dS_{3}$ static patch as explained in Coleman:2021nor . It does not capture local bulk matter details treated by a more elaborate algorithm in Batra:2024kjl .
@@
which will lead to a simple differential equation for the energy spectrum. As just noted, we will denote the solvable $T^{2}$ deformation with the replacement (5) by $(T^{2})_{s}$. The expressions (5) faithfully capture only the entropically dominant bands of energy in the system, but the differential equation (4) with this substitution can be solved for all energy levels. This is analogous to the situation in $d=2$, where the pure $T\bar{T}+\Lambda_{2}$ theory accurately captures the most entropic energy bands, but it can be applied to all energy levels. In that case, accurately capturing general matter excitations requires an additional prescription Batra:2024kjl . In fact in contrast to the $d=2$ case, we do not faithfully capture angular momentum states in the solvable theory that we obtain using (5). For the $d=3$ (and higher) case of interest in this work, we relegate to the additional algorithmic prescription in §4 the faithful description of angular momentum states. However, it might be possible to devise a more general solvable theory that captures faithfully the angular momentum states (rather than just the entropically dominant states). For more discussion of the possibility of extending the solvable theory to capture them see the end of §3 below.
@@
We then increase $y$ with the lower (positive) sign of the square root. This builds a model reproducing the horizon entropy – the dimension of the finite real Hilbert space – and radial geometry of the bounded static patch for the $dS_{4}$ model. The horizon entropy includes the leading Gibbons-Hawking Gibbons:1977mu result $A/4g_{N}$ along with any universal logarithmic corrections $\propto\log(A/4G_{N})$ tied to the symmetry structure in the bounded AdS and dS geometries Anninos:2020hfj ; Chandrasekaran:2022cip . The radial geometry is inferred from the map between the deformed energy as a function of $y$ and the Brown-York energy as a function of the size of the bounding cylinder on the gravity side, for the appropriate band of energies corresponding to the gravity-side geometry.
@@ 4.1 The uplift part of the deformation
In this section we will outline the steps in the algorithm which resolve the uplift from negative to positive bulk cosmological constant $\Lambda_{4}\propto-\eta/\ell_{AdS}^{2}$. At the level of the solvable model, this amounts to changing $\eta$. In the full algorithm, we use fine tuning to accurately treat the matter sectors which effect this uplift. These include a scalar $\Phi_{u}$ which interpolates from an AdS minimum (such as the M2-brane CFT Maldacena:1997re $AdS_{4}\times S^{7}$ with $F_{7}$ flux in M theory) to a dS minimum (such as a member of the discretuum in the explicit compactifications DeLuca:2021pej ). In the examples DeLuca:2021pej , this interpolation involves a topology change; as mentioned earlier this assumes that the M theory configuration space is connected, similarly to stringy topology changing transitions studied in Adams:2005rb .
@@
The spectrum of light EFT fields in the bulk changes during the uplift. In general, the bulk action $S_{bulk}$ contains interactions between $\Phi_{u}$ and other fields, which we will collectively denote $\Phi_{\perp}$. In particular, the potential $V(\Phi_{u},\Phi_{\perp})$ will contain $\Phi_{u}$-dependent mass terms for the $\Phi_{\perp}$ fields.
@@
The uplift consists of changing the boundary condition for $\Phi_{u}$, denoted $\Phi_{uc}$, from its AdS to its dS value. In the boundary theory variables, this corresponds to changing the corresponding coupling constant in the theory. We wish to derive the analogue of (40) for this uplift process, determining the effect on the spectrum of this interpolation. We apply the standard bulk/boundary dictionary as laid out in Gubser:1998bc ; Witten:1998qj and Balasubramanian:1999re . Namely, at the level of large-N saddle point solutions and their corresponding energy bands in the boundary theory, we have
$\partial_{\Phi_{cu}}\langle n|H|n\rangle_{GR+EFT+\dots}=-\partial_{\Phi_{uc}}S_{bulk}\big{[}{\bf\Phi}_{*,n}(\Phi_{uc})\big{]}$ (43)
@@
where ${\bf\Phi}_{*}(\Phi_{uc})$ denotes the $nth$ saddle point solution for the $\Phi_{u},\Phi_{\perp}$ fields, which depends on the coupling $\Phi_{uc}$. As in the discussion above around (41), the fine grained states are matched to the closest energy state among this cruder state space resolvable in $GR+EFT+\dots$, up to negligible differences that can be included without affecting the approximate agreement to the bulk. This leads to blocks of nearly degenerate states in our finite Hamiltonian. The total number of states $\exp(S)$ contains logarithmic corrections to the entropy. This includes certain corrections $\propto\log(A/4G_{N})$ to the entropy which follows from the symmetry structure in our bounded geometry (as explained in Anninos:2020hfj ; Chandrasekaran:2022cip ). [^11] We thank Dio Anninos and Albert Law for very helpful discussions of this point. Since the symmetry structure is the same in the bounded AdS black hole and the bounded dS cosmic horizon patch, the result for this contribution should automatically match as we proceed through the uplift. Determining this coefficient would arise from a calculation analogous to those in Anninos:2020hfj ; Chandrasekaran:2022cip and Sen:2012dw . [^12] It may also be possible to generalize the CFT analysis of Benjamin:2023qsc to our case. It would be interesting to analyze this point further, in order to determine whether the matching of the full EFT + GR +…prediction for the corrected entropy is automatic, including matter contributions as well as gravitational and matter counterterms Anninos:2020hfj . If not, it would be possible to make small adjustments to the state count in our spectrum to match the bulk effective theory prediction.

### v3

@@ 1 Introduction
The results Coleman:2021nor ; Batra:2024kjl were specific to a 3-dimensional bulk theory. In this note, we refine the method and point out that it readily generalizes to the case of four bulk dimensions, even though the special properties of the $T\bar{T}$ operator in two dimensions Zamolodchikov:2004ce ; Smirnov:2016lqw do not extend to the full $T^{2}$ operator in higher dimensions Hartman:2018tkw ; Taylor:2018xcy . For boundary dimension $d=3$ there is still a sector – containing the most entropic bands of energy levels – that is captured by a solvable deformation of the $AdS_{4}/CFT_{3}$ Hamiltonian similar to Coleman:2021nor . We define this solvable theory by restricting the $T^{2}$ deformation of AdS/CFT (as derived at large $N$ in Hartman:2018tkw and generalized here to include a constant $\Lambda_{3}$ addition at each step) to contain only diagonal components of the stress tensor and to take the form that that it does in symmetric homogeneous states, depending only on energy density $E/V$ and pressure $-dE/dV$. The finite real spectrum of this Hamiltonian, manifest in equation (6) below, captures the Gibbons-Hawking entropy along with the radial geometry of the $dS_{4}$ cosmic horizon patch as a function of boundary size. The operator algebra, working at finite $N$, is type I, simpler than the type II result one obtains in the $G_{N}\to 0$ limit Chandrasekaran:2022cip .
@@ 1.1 Status of classical and quantum timelike boundaries
From the classical, low energy perspective it is well known that the initial boundary value problem is not well defined for broad classes of Dirichlet boundary conditions, at least linearized around simple backgrounds. For example, the radial constraint linearized around these backgrounds cannot be satisfied for an arbitrary fixed boundary geometry An:2021fcq . This has been straightforwardly proved for perturbations of a flat boundary in Minkowski spacetime (which we will refer to as a ‘flat corner’) to nearby boundary metrics with nonzero intrinsic scalar curvature $\delta{\mathcal{R}}$: with vanishing unperturbed extrinsic curvature, the linearized perturbation to the radial constraint simply reads $\delta{\mathcal{R}}=0$. (This argument does not exclude solutions of the radial constraint with nontrivial extrinsic curvature scaling like $K\sim\sqrt{\delta g}$, which exists in many cases. Still, according to an argument in terms of Sobolev spaces in Anderson:2010ph these nonlinear solutions do not exist for generic fixed boundary metrics, as was seen in early works on near horizon physics Bredberg:2011xw ; Anninos:2011zn .) At least for flat corners at the linearized level there is also a violation of geometric uniqueness Anninos:2022ujl ; An:2021fcq : given initial data, it does not evolve to a unique solution at later times. At least for the Euclidean case, boundaries with zero extrinsic curvature also lead to problems defining a propagator for perturbation theory Anderson:2006lqb ; Witten:2018lgb .
@@ 2 A solvable $T^{2}+\Lambda_{3}$ deformation at finite N
In the prescription Coleman:2021nor , one starts with $\eta=1$, deforms until $y$ reaches $y_{0}=3/c\pi^{2}$, interpolates to $\eta=-1$ (resolved into an uplift achieved by appropriate matter Dong:2010pm in Batra:2024kjl ), and proceeds back toward larger $y$ with the $\Lambda_{2}$ terms in place. For this choice of $y_{0}$, the top band of energy levels corresponds to the boundary sitting at the horizon of the $\Delta=c/6$ BTZ black hole. This maps onto the empty dS cosmic patch upon continuing with the $T\bar{T}+\Lambda_{2}$ deformation. This by itself accounts for the refined Gibbons-Hawking entropy and radial geometry of the bounded $dS_{3}$ static patch as explained in Coleman:2021nor . It does not capture local bulk matter details treated by a more elaborate algorithm in Batra:2024kjl .
@@
which will lead to a simple differential equation for the energy spectrum. As just noted, we will denote the solvable $T^{2}$ deformation with the replacement (5) by $(T^{2})_{s}$, which is not a local deformation. The expressions (5) faithfully capture only the entropically dominant bands of energy in the system, but the differential equation (4) with this substitution can be solved for all energy levels. In this respect, it is analogous to the situation in $d=2$, where the pure $T\bar{T}+\Lambda_{2}$ theory accurately captures the most entropic energy bands, but it can be applied to all energy levels. In that case, accurately capturing general matter excitations requires an additional prescription Batra:2024kjl . In fact in contrast to the $d=2$ case, we do not faithfully capture angular momentum states in the solvable theory that we obtain using (5). For the $d=3$ (and higher) case of interest in this work, we relegate to the additional algorithmic prescription in §4 the faithful description of angular momentum states. However, it might be possible to devise a more general solvable theory that captures faithfully the angular momentum states (rather than just the entropically dominant states). For more discussion of the possibility of extending the solvable theory to capture them see the end of §3 below.
@@
The previous literature Coleman:2021nor Batra:2024kjl focused for simplicity on $\eta_{u}=-1$, but given the fine discretuum of de Sitter models DeLuca:2021pej we can readily generalize away from that case. We then increase $y$ with the lower (positive) sign of the square root. This builds a model reproducing the horizon entropy – the dimension of the finite real Hilbert space – and radial geometry of the bounded static patch for the $dS_{4}$ model. The horizon entropy includes the leading Gibbons-Hawking Gibbons:1977mu result $A/4g_{N}$ along with any universal logarithmic corrections $\propto\log(A/4G_{N})$ tied to the symmetry structure in the bounded AdS and dS geometries Anninos:2020hfj ; Chandrasekaran:2022cip . The radial geometry is inferred from the map between the deformed energy as a function of $y$ and the Brown-York energy as a function of the size of the bounding cylinder on the gravity side, for the appropriate band of energies corresponding to the gravity-side geometry.
@@
It may be useful to clarify that, while in $d=2$ the BH temperature measured on the AdS boundary and that associated to the static patch happen to be the same, they differ for $d>2$. However, our procedure involves matching microcanonical parameters –energies and volumes– at the horizons, and the difference between the temperature in the asymptotic AdS region and that for timelike-bounded dS at the pole are not directly relevant, because the temperature (or ‘tomperature’ Lin:2022nss) evolves along our trajectory. This evolution of the temperatures along the trajectory will also be reflected in the matter bulk dynamics, such as the periodicities of 2-point functions. The fine-tuning procedure of Sec. 4 will account for these properties.
@@ 4.1 The uplift part of the deformation
In this section we will outline the steps in the algorithm which resolve the uplift from negative to positive bulk cosmological constant $\Lambda_{4}\propto-\eta/\ell_{AdS}^{2}$. At the level of the solvable model, this amounts to changing $\eta$. In the full algorithm, we use fine tuning to accurately treat the matter sectors which effect this uplift. These include a scalar $\Phi_{u}$ which interpolates from an AdS minimum (such as the M2-brane CFT Maldacena:1997re $AdS_{4}\times S^{7}$ with $F_{7}$ flux in M theory) to a dS minimum (such as a member of the discretuum in the explicit compactifications DeLuca:2021pej ). In the examples DeLuca:2021pej , this interpolation involves a topology change; as mentioned earlier this assumes that the M theory configuration space is connected, similarly to stringy topology changing transitions studied in Adams:2005rb , with $\Phi_{u}$ representing the condensate effecting the topology change analagous to the winding modes in Adams:2005rb .
@@
The spectrum of light EFT fields in the bulk changes during the uplift. In general, the bulk action $S_{bulk}$ contains interactions between $\Phi_{u}$ and other fields, which we will collectively denote $\Phi_{\perp}$. In particular, the potential $V(\Phi_{u},\Phi_{\perp})$ will contain $\Phi_{u}$-dependent mass terms for the $\Phi_{\perp}$ fields, as in the related discussion in Batra:2024kjl .
@@
The uplift consists of changing the boundary condition for $\Phi_{u}$, denoted $\Phi_{uc}$, from its AdS to its dS value. In the boundary theory variables, this corresponds to changing the corresponding coupling constant $g_{u}$ in the theory. We wish to derive the analogue of (40) for this uplift process, determining the effect on the spectrum of this interpolation. We apply the standard bulk/boundary dictionary as laid out in Gubser:1998bc ; Witten:1998qj and Balasubramanian:1999re . Namely, at the level of large-N saddle point solutions and their corresponding energy bands in the boundary theory, we have
$\partial_{g_{u}}\langle n|H|n\rangle_{def-CFT}\equiv\partial_{\Phi_{cu}}\langle n|H|n\rangle_{GR+EFT+\dots}=-\partial_{\Phi_{uc}}S_{bulk}\big{[}{\bf\Phi}_{*,n}(\Phi_{uc})\big{]}$ (43)
@@
where ${\bf\Phi}_{*}(\Phi_{uc})$ denotes the $nth$ saddle point solution for the $\Phi_{u},\Phi_{\perp}$ fields, which depends on the coupling $\Phi_{uc}$. As in the discussion above around (41), the fine grained states are matched to the closest energy state among this cruder state space resolvable in $GR+EFT+\dots$, up to negligible differences that can be included without affecting the approximate agreement to the bulk. That is, $|n\rangle_{GR+EFT+\dots}$ is the state in the resolvable spectrum of $GR+EFT+\dots$ with energy that is closest to that of $|n\rangle_{def-CFT}$ in the deformed-CFT spectrum. This leads to blocks of nearly degenerate states in our finite Hamiltonian. The total number of states $\exp(S)$ contains logarithmic corrections to the entropy. This includes certain corrections $\propto\log(A/4G_{N})$ to the entropy which follows from the symmetry structure in our bounded geometry (as explained in Anninos:2020hfj ; Chandrasekaran:2022cip ). [^11] We thank Dio Anninos and Albert Law for very helpful discussions of this point. Since the symmetry structure is the same in the bounded AdS black hole and the bounded dS cosmic horizon patch, the result for this contribution should automatically match as we proceed through the uplift. Determining this coefficient would arise from a calculation analogous to those in Anninos:2020hfj ; Chandrasekaran:2022cip and Sen:2012dw . [^12] It may also be possible to generalize the CFT analysis of Benjamin:2023qsc to our case. It would be interesting to analyze this point further, in order to determine whether the matching of the full EFT + GR +…prediction for the corrected entropy is automatic, including matter contributions as well as gravitational and matter counterterms Anninos:2020hfj . If not, it would be possible to make small adjustments to the state count in our spectrum to match the bulk effective theory prediction.